OpenStudy (anonymous):

have school in a little and forgot to do a few problems someone please help, emergency ! 3x+y=14 x=y-3

OpenStudy (turingtest):

substitute x=y-3 into the first equation to get 3(y-3)+y=14 and solve for y

OpenStudy (turingtest):

the second equation will then give you x

OpenStudy (anonymous):

substitution or elimination you can use either of the two

OpenStudy (anonymous):

3(y-3)+y = 14 3y - 9 + y = 14 4y - 9 = 14 4y = 14+9 4y = 23 y = 23/4 = 5.75 Substituting y = 5.75 : x= y - 3 x = (5.75) - 3 = 2.75